# Unfolding AlphaFold’s Bayesian Roots in Probability Kinematics

### v2

@@ 5 Experiments
where $B(\alpha,\beta)$ is the standard normalization constant of the Beta distribution. We will write $r\sim\textrm{ScaledBeta}(r;\alpha,\beta,N)$. In order to apply PK, we assume by construction that the J-condition, $p(\boldsymbol{\theta}\mid r)=\pi(\boldsymbol{\theta}\mid r)$, holds. That is, prior $\pi_{\textrm{VRW}}(\boldsymbol{\theta})$ and posterior $p(\boldsymbol{\theta})$ share the same marginal with respect to $r$. We can apply PK to update the prior, following $p(\boldsymbol{\theta})=\int\pi_{\textrm{VRW}}(\boldsymbol{\theta}\mid r)p(r)\textrm{d}\boldsymbol{\theta}$, but the marginal $\pi_{\textrm{VRW}}(\boldsymbol{\theta}\mid r)$ is not available in tractable form for a VRW. Therefore, as is the case for AF1, we use the RR update given by (14), making use of the fact that the marginal over $r$ according to the prior is given by (22). The PK update has the following structure,
$p(\boldsymbol{\theta})$ $=\frac{p\left(r(\boldsymbol{\theta})\right)}{\pi_{\textrm{VRW}}\left(r(\boldsymbol{\theta})\right)}\pi_{\textrm{VRW}}(\boldsymbol{\theta}),$ (25)
@@
which results in the final model,
@@
$p(\boldsymbol{\theta};\mu,\kappa,\alpha,\beta,N)$ $=\frac{\textrm{ScaledBeta}\left(r(\boldsymbol{\theta});\alpha,\beta,N\right)}{\textrm{Stephens}\left(r(\boldsymbol{\theta});\kappa,N\right)}\pi_{\textrm{VRW}}(\boldsymbol{\theta};\mu,\kappa,N).$ (26)
@@
We sample from this model using the No U-Turn Sampler (NUTS) [29], as implemented in in the deep probabilistic programming language Pyro [6], with default parameters. We use 1000 warm up and 1000 posterior samples, taking under one minute on a standard laptop with an Intel i7 CPU.
@@
We set $N=5$ and $\mu=0,\kappa=10$ for the VRW, and $\textrm{ScaledBeta}(\alpha=10,\beta=10,N=5)$ as target. Figure 3 visually confirms that the correct posterior is attained. To further evaluate the model, we compared the empirical distributions of $r$ to the target distribution using the one-sample Kolmogorov–Smirnov (KS) test [68], including every fifth NUTS sample to reduce autocorrelation. Across 10 NUTS simulations, the KS statistic ranges from $D=0.04$ to $D=0.08$, with a corresponding median p-value of $0.41$, indicating a good match.
@@
In contrast and as expected, a naive reweighing approach that omits the reference distribution produced a much poorer fit, with the KS statistic ranging from $D=0.62$ to $D=0.70$, and $p<10^{-6}$ in all cases. The deviation is also very obvious visually (Figure 3, gray histogram). This ablation experiment confirms the importance of including the reference distribution.
6 Discussion
KBPs [58] were quantitatively explained by appealing to (14) by Hamelryck et al. [25, 47, 45, 67], and later recognized as an instance of PK [27]. However, these references predate AF1, and are unconnected to deep learning. Moreover, the connection between KBPs and PK, while formally sound, has remained virtually unexplored within the deep learning paradigm applied to biomolecular structure. To our knowledge, our work is the first to explain AF1 as an instance of PK.
Although PK is a well-defined concept in probability theory [32, 16, 33, 13, 48], as far as we know, it has seen almost no explicit adoption in deep models. Yu (2020) applies PK to training Bayesian neural networks using soft evidence [70]. More recently, Munk et al. (2023) applied PK to the incorporation of soft evidence in probabilistic models and stochastic simulators [50].
This work connects physics-inspired modeling, probabilistic reasoning, and deep learning, which are increasingly converging [18]. The success of AF1 and its roots in PK hint at a general approach to formulate deep models that are compositional [22, 6, 69]. This approach consists of updating an empirical Bayes prior concerning a fine-grained variable with soft evidence on a coarse-grained variable, both obtained from deep models. We point out that PK is agnostic regarding the nature of its components, bringing a wide range of deep generative models potentially within its scope [66].
@@ 7 Conclusions
We reinterpreted AF1 as an instance of PK rather than as a thermodynamic potential, and demonstrated the probabilistic framework using a synthetic model that allows qualitative evaluation. Beyond clarifying the foundations of a paradigm-shifting method, our reinterpretation and AF1’s well-established accomplishments imply that PK offers a principled, quantative foundation for building compositional deep probabilistic models based on generalized Bayesian updating [22, 6, 69].
@@ 8 Acknowledgments
The authors thank Wouter Boomsma, Jesper Ferkinghoff-Borg, Jes Frellsen, Carsten Hartmann, John Kent, Ola Rønning, Manfred Sippl and Sandy Zabell for discussions and comments, and acknowledge funding from the VILLUM Experiment Programme (grant 50240).
@@
9 Example: Whitworth’s horses
This simple application of PK is discussed by Diaconis and Zabell [16]. Three horses, A, B, and C, have prior winning probabilities $\pi(A)=1/2$, $\pi(B)=1/4$, and $\pi(C)=1/4$, respectively. New evidence increases the probability of A winning from $\pi(A)=1/2$ to $p(A)=2/3$, without providing any further information on the other horses. Standard Bayesian conditioning cannot apply, as no specific event has occurred. Instead, PK prescribes an update that satisfies the new marginal, $p(A)=2/3$, while preserving the conditional probabilities, $p(B\mid A)=p(C\mid A)=1/2$. This yields the updated posterior probabilities, $p(A)=2/3$, $p(\sim A)=1/3$, and
$p(B)=p(C)=p(B\mid\sim A)p(\sim A)=p(C\mid\sim A)p(\sim A)=1/2*1/3=1/6,$ (27)
@@ 9 Example: Whitworth’s horses
10 Local nature of the prior and J-condition
@@
As the dihedral angles are sufficient to generate 3D structures, one might ask why it is necessary to add evidence on pairwise distances. It is well known that probabilistic models based on internal coordinates, such as dihedral angles, struggle to reproduce realistic distances between amino acids that are far apart in sequence position [4]. This is due to the highly structured covariance induced by the chain geometry: small local perturbations in dihedral angles propagate in a correlated way through the backbone.
@@
To illustrate this in the context of AF1, we sampled dihedral angles from the prior and reconstructed the corresponding 3D coordinates. AF1 code that can be executed for CASP13 targets is available from GitHub. [^3] https://github.com/google-deepmind/deepmind-research/tree/master/alphafold_casp13 AF1 was applied to CASP13 target TS1019s2, corresponding to the PT-VENN domain-containing protein, Protein Data Bank Identifier 8EY4, chain B. For each amino acid, the prior is specified as probabilities over the Ramachandran plot, binned at $10^{\circ}\times 10^{\circ}$ resolution, obtained from a deep model (see Section 4) [56]. For each amino acid, we picked a random bin according to the specified probabilities, and set the dihedral angles to the bin’s centroid. We then construct the 3D coordinates of the $N,C\alpha,C$ backbone from the sampled dihedral angles. The sampled structures were superimposed on the true structure using a standard algorithm [35]. Both prior probabilities and sampled 3D structures are available as supplementary material in the ZIP file (see README file).
Figure 4: Distogram of the native structure (left) and the averaged distogram over 100 samples from the AF1 prior (right) for the PT-VENN domain-containing protein (Protein Data Bank Identifier 8EY4, Chain B). Distances are in Ångstrøm.
Figure 5: 100 samples from the AF1 prior (transparent, thin ribbons) superimposed on the cartoon representation of the native structure of the PT-VENN domain-containing protein (Protein Data Bank Identifier 8EY4, Chain B). In all cases, the N-terminus is shown in blue and C-terminus in red. Figure made with PyMol [54].
@@
11 Practical aspects of the AF1 potential
Our statistical framework explains AF1’s potential as the update of a prior over dihedral angles with information on pairwise distances. For practical reasons, the application of PK by AF1 is heuristic in a few respects, which we clarify here. First, the reference distribution over the coarse grained variable, $\pi(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi});\boldsymbol{\varGamma})$, is not directly obtained from the prior over the fine-grained variable, $\pi(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta})$. Rather, it is approximated by training a model that uses as input only the sequence length, $L$, and vector specifying the position of the Glycine residues.
Second, AF1 does not use proper PDFs over the distance matrix, $\boldsymbol{D}$. For $n$ points in $\mathbb{R}^{d}$, the n(n1)/2n(n-1)/2 pairwise distances cannot vary independently: they must satisfy geometric constraints to correspond to an embeddable configuration. As discussed by Clifford and Green (1985) [15], valid distances lie on a smaller manifold of dimension ndd(d+1)/2nd-d(d+1)/2. By modeling probability distribution over $\boldsymbol{D}$ as a product of independent marginals, AF1 defines an approximate distribution that is singular, i.e., it may place nonzero density on geometrically invalid configurations.

### v3

@@ 5 EXPERIMENTS
We can apply PK to update the prior, following $p(\boldsymbol{\theta})=\int\pi_{\textrm{VRW}}(\boldsymbol{\theta}\mid d)p(d)\textrm{d}\boldsymbol{\theta}$, but the marginal $\pi_{\textrm{VRW}}(\boldsymbol{\theta}\mid d)$ is not available in a tractable form for a VRW. Therefore, as is the case for AF1, we use PK’s reference ratio update given by (16), making use of the fact that the marginal over $d$ according to the prior is given by (27). The PK update has the following form,
$p(\boldsymbol{\theta})$ $=\frac{p\left(d(\boldsymbol{\theta})\right)}{\pi_{\textrm{VRW}}\left(d(\boldsymbol{\theta})\right)}\pi_{\textrm{VRW}}(\boldsymbol{\theta}),$ (30)
@@
which, after filling in the concrete PDFs, results in the final model,
@@
$p(\boldsymbol{\theta};\mu,\kappa,\alpha,\beta,N)=\\ \frac{\textrm{ScaledBeta}\left(d(\boldsymbol{\theta});\alpha,\beta,N\right)}{\textrm{Stephens}\left(d(\boldsymbol{\theta});\kappa,N\right)}\pi_{\textrm{VRW}}(\boldsymbol{\theta};\mu,\kappa,N).$ (31)
@@
We sample from this model using the No-U-Turn Sampler (NUTS) (Hoffman and Gelman, 2014), as implemented in the deep probabilistic programming language Pyro (Bingham et al., 2019), with default parameters. We use 1000 warm-up and 1000 posterior samples, taking under one minute on a standard laptop with an Intel i7 CPU. For code availability, see Appendix 8.
@@
We set $N=5$ and $\mu=0,\kappa=10$ for the VRW, and $\textrm{ScaledBeta}(\alpha=10,\beta=10,N=5)$ as target. Figure 3 visually confirms that the correct posterior is attained. To further evaluate the model, we compared the empirical distributions of $d$ to the target distribution using the one-sample Kolmogorov–Smirnov (KS) test (Virtanen et al., 2020), including every fifth NUTS sample to reduce autocorrelation. Across 10 NUTS simulations, the KS statistic ranges from $0.04$ to $0.08$, with a corresponding median p-value of $0.41$, indicating a good match.
@@
In contrast and as expected, a naive reweighting approach that omits the reference distribution produced a much poorer fit, with the KS statistic ranging from $0.62$ to $0.70$, and $p<10^{-70}$ in all cases. The deviation is also very obvious visually (Figure 3, gray histogram). This ablation experiment confirms the importance of including the reference distribution.
6 DISCUSSION AND RELATED WORK
KBPs (Sippl, 1990) were quantitatively explained by appealing to (16) by Hamelryck et al. (Hamelryck et al., 2010; Mardia et al., 2011; Mardia and Hamelryck, 2012; Valentin et al., 2014), and later recognized as an instance of PK (Hamelryck et al., 2015). However, these references predate AF1, and are unconnected to deep learning. Moreover, the connection between KBPs and PK, while formally sound, has remained virtually unexplored within the deep learning paradigm applied to biomolecular structure. To our knowledge, our work is the first to analyze AF1 as an instance of PK.
Although PK is a well-defined concept in probability theory (Jeffrey, 1990; Diaconis and Zabell, 1982; Jeffrey, 2004; Chan and Darwiche, 2005; Meehan and Zhang, 2020), as far as we know, it has seen almost no explicit adoption in deep models: Yu (2020) applies PK to training Bayesian neural networks with soft evidence. Munk et al. (2023) use PK to incorporate soft evidence in probabilistic models and stochastic simulators.
This work connects statistical mechanics, probabilistic modeling, and deep learning, which are increasingly converging in the field of protein structure prediction (Durumeric et al., 2023; Jing et al., 2025; Olsson, 2026). The success of AF1 and its roots in PK hint at a general approach to formulate deep models that are compositional (Ghahramani, 2015; Bingham et al., 2019; Wood et al., 2014). This approach consists of updating an empirical Bayes prior concerning a fine-grained variable with soft evidence on a coarse-grained variable, both obtained from deep models. We point out that PK is agnostic regarding the nature of its components, bringing a wide range of deep generative models potentially within its scope (Tomczak, 2024).
Finally, probability kinematics has been connected to credal set theory, where uncertainty is represented by closed, convex sets of probability distributions rather than a single posterior (Marchetti and Antonucci, 2018; Caprio and Gong, 2023; Caprio and Mukherjee, 2023), providing opportunities for formulating deep generative models that can explicitly accommodate imprecise, partially conflicting or even inconsistent information.
@@ 7 CONCLUSIONS
We interpret AF1 as an instance of PK and demonstrated the probabilistic framework using a synthetic model with qualitative evaluation. Beyond clarifying the foundations of a paradigm-shifting method, our results imply that PK offers a principled foundation for building compositional deep probabilistic models based on generalized Bayesian updating (Ghahramani, 2015; Bingham et al., 2019; Wood et al., 2014).
8 CODE AVAILABILITY
Code for reproducing the synthetic von Mises random walk experiment in Section 5 is available at: https://github.com/thamelry/AlphaFoldPK.
@@ 9 ACKNOWLEDGMENTS
We thank Wouter Boomsma, Jesper Ferkinghoff-Borg, Jes Frellsen, Carsten Hartmann, John T. Kent, Ola Rønning, Manfred Sippl and Sandy Zabell for discussions and comments, and an anonymous reviewer for pointing out the connection to credal sets. We acknowledge funding from the VILLUM Experiment Programme (grant 50240) and the Independent Research Fund Denmark (grant 10.46540/5244-00031B).
@@
Appendix A NOTATION AND ABBREVIATIONS
Symbol / Abbreviation Meaning (and shape of arrays)
AF1 AlphaFold version 1
KBP Knowledge Based Potential
MAP Maximum A Posteriori estimate
MSA Multiple Sequence Alignment
NN Neural Network
$\mathcal{N}(\mu,\sigma^{2})$ Normal distribution with mean $\mu$ and variance $\sigma^{2}$
PK Probability Kinematics (Jeffrey conditioning)
PMF Potential of Mean Force
[…]
VRW von Mises Random Walk
$\pi(\cdot)$ Prior PDF
AF1
$L$ Length (number of amino acids) of a protein
$\boldsymbol{s}$ Protein sequence, one-hot encoding, shape $L\times 20$
$\boldsymbol{M}$ MSA, shape $M\times L\times 20$ for $M$ sequences
$\mathbb{I}_{\textrm{Gly}}(\boldsymbol{s})$ Glycines in protein sequence, one-hot encoding, shape $L$
$\boldsymbol{X}$ Coordinates of protein backbone, shape $L\times 4\times 3$
$\boldsymbol{X}(\boldsymbol{\phi},\boldsymbol{\psi})$ Realization of $\boldsymbol{X}$ for given dihedral angles
$\boldsymbol{\phi},\boldsymbol{\psi}$ Protein dihedral angles in $(-\pi,\pi]$, shape $L$ each
$\boldsymbol{D}$ $\beta-C$$\beta$ distance matrix, shape $L\times L$
$\boldsymbol{D}(\boldsymbol{X}),\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi})$ Realization of $\boldsymbol{D}$ for given coordinates or angles
$\pi(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta})$ Prior over dihedral angles (von Mises)
$\pi(\boldsymbol{D};\boldsymbol{\Gamma})$ Prior over pairwise distances
$p(\boldsymbol{D};\boldsymbol{\Lambda})$ Target PDF of $\boldsymbol{D}$, used for PK update
Synthetic model (von Mises random walk, VRW)
NN Number of random walk steps
$\boldsymbol{\chi}$ 2D coordinates of VRW starting at $(0,0)$, shape $(N+1)\times 2$
$\boldsymbol{\theta}$ Angles of a VRW, shape NN
$d$ Origin-to-end distance of a VRW, $d\in\mathbb{R}^{+}$
$d(\boldsymbol{\chi}),d(\boldsymbol{\theta})$ Realization of $d$ for given VRW 2D coordinates or angles
$\pi_{\textrm{VRW}}(\boldsymbol{\theta};\mu,\kappa,N)$ Prior over VRW angles (von Mises)
[…]
PMF and PK summary
$\boldsymbol{\omega}$ A fine-grained random variable (e.g., $\boldsymbol{X}$)
$\boldsymbol{\xi}$ A coarse-grained or collective random variable (e.g., $\boldsymbol{D}$)
$\boldsymbol{\xi}(\boldsymbol{\omega})$ Realization of $\boldsymbol{\xi}$ given $\boldsymbol{\omega}$ (e.g., $\boldsymbol{D}(\boldsymbol{X})$)
J-condition Jeffrey condition, posterior $p(\boldsymbol{\omega}\mid\boldsymbol{\xi})=\pi(\boldsymbol{\omega}\mid\boldsymbol{\xi})$
PMF; maps $p(\boldsymbol{\omega})$ to $p(\boldsymbol{\xi})$ $p(\boldsymbol{\xi})=\int\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))p(\boldsymbol{\omega})\textrm{d}\boldsymbol{\omega}$
PK; updates $\pi(\boldsymbol{\omega})$ with $p(\boldsymbol{\xi})$ $p(\boldsymbol{\omega})=\int\pi(\boldsymbol{\omega}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})\textrm{d}\boldsymbol{\xi}$
Reference ratio formulation of PK $p(\boldsymbol{\omega})=\int\frac{\delta(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{\omega}))\pi(\boldsymbol{\omega})}{\pi(\boldsymbol{\xi})}p(\boldsymbol{\xi})\textrm{d}\boldsymbol{\xi}=\frac{p(\boldsymbol{\xi}(\boldsymbol{\omega}))}{\pi(\boldsymbol{\xi}(\boldsymbol{\omega}))}\pi(\boldsymbol{\omega})$
Appendix B AF1 AND THE SYNTHETIC MODEL
Table 1 summarizes the conceptual correspondences between AF1 and our synthetic von Mises random walk (VRW) model. Both models share the same probabilistic structure: a prior distribution over (fine-grained) angular variables induces Euclidean coordinates, which in turn determine a (coarse-grained) distance variable. The map from angles to distances is many-to-one in both cases. In AF1, dihedral angles generate 3D coordinates, $\boldsymbol{X}$, and corresponding $C\beta-C\beta$ distances; in the synthetic model, VRW angles generate 2D coordinates and an origin-to-end distance or resultant length, $d$. This analogy ensures that the PK update plays the same role in both cases, i.e., an angular prior is updated with distance information, allowing us to study the probabilistic foundation of AF1 in a tractable and interpretable setting.
Table 1: Comparison of AF1 and the synthetic model.
AF1 Synthetic model Correspondence
Dihedral angles, ($\boldsymbol{\phi},\boldsymbol{\psi}$) VRW angles, $\boldsymbol{\theta}$ Prior over angles (von Mises)
3D atomic coordinates, $\boldsymbol{X}$ 2D VRW coordinates, $\boldsymbol{\chi}$ Coordinates generated by angles
Pairwise distances, $\boldsymbol{D}$ Resultant length, $d$ PK update with distance information
Appendix C WHITWORTH’S HORSES
This simple application of PK is discussed by Diaconis and Zabell (Diaconis and Zabell, 1982). Three horses, A, B, and C, have prior winning probabilities $\pi(A)=1/2$, $\pi(B)=1/4$, and $\pi(C)=1/4$, respectively. New evidence increases the probability of A winning from $\pi(A)=1/2$ to $p(A)=2/3$, without providing any further information on the other horses. Standard Bayesian conditioning cannot apply, as no specific event has occurred. Instead, PK prescribes an update that satisfies the new marginal, $p(A)=2/3$ and $p(A^{c})=1/3$, while preserving the conditional probabilities, $\pi(B\mid A^{c})=\pi(C\mid A^{c})=1/2$. This yields the updated posterior probabilities,
$p(B)$ $=\pi(B\mid A^{c})p(A^{c})=1/2\cdot 1/3=1/6,$ (32)
$p(C)$ $=\pi(C\mid A^{c})p(A^{c})=1/2\cdot 1/3=1/6,$ (33)
@@ Appendix C WHITWORTH’S HORSES
Appendix D LOCAL NATURE OF AF1 PRIOR AND J-CONDITION
@@
As the dihedral angles are sufficient to generate 3D structures, one might ask why it is necessary to add evidence on pairwise distances. It is well known that probabilistic models based on internal coordinates, such as dihedral angles, struggle to reproduce realistic distances between amino acids that are far apart in sequence position (Arts et al., 2024). This is due to the highly structured covariance induced by the chain geometry: small local perturbations in dihedral angles propagate in a correlated way through the backbone.
@@
To illustrate this in the context of AF1, we sampled dihedral angles from the prior and reconstructed the corresponding 3D coordinates. AF1 code that can be executed for CASP13 targets is available from GitHub. [^5] https://github.com/google-deepmind/deepmind-research/tree/master/alphafold_casp13 AF1 was applied to CASP13 target TS1019s2, corresponding to the PT-VENN domain-containing protein, Protein Data Bank Identifier 8EY4, chain B. For each amino acid, the prior is specified as probabilities over the Ramachandran plot, binned at $10^{\circ}\times 10^{\circ}$ resolution, obtained from a deep model (see Section 4) Senior et al. (2020). For each amino acid, we picked a random bin according to the specified probabilities, and set the dihedral angles to the bin’s centroid. We then construct the 3D coordinates of the $N,C\alpha,C$ backbone from the sampled dihedral angles. The sampled structures were superimposed on the true structure using a standard algorithm Kabsch (1976); Dryden and Mardia (2016).
Figure 4: Distogram of the native structure (left) and the averaged distogram over 100 samples from the AF1 prior (right) for the PT-VENN domain-containing protein (Protein Data Bank Identifier 8EY4, Chain B). Distances are in Ångström. The prior accurately captures short-range distances between amino acids neighboring in sequence (i.e., distances close to the diagonal in the distogram on the right), but not long-range distances between amino acids far apart in the sequence, which necessitates updating the prior with distance information.
Figure 5: 100 samples from the AF1 prior (transparent, thin ribbons) superimposed on the cartoon representation of the native structure of the PT-VENN domain-containing protein (Protein Data Bank Identifier 8EY4, Chain B). In all cases, the N-terminus is shown in blue and C-terminus in red. Note how the samples from the prior are spatially extended compared to the compact native structure, which is representative for probabilistic models based on dihedral angles (Arts et al., 2024). Figure made with PyMol Schrödinger, LLC (2015).
@@
Appendix E PRACTICAL ASPECTS OF THE AF1 POTENTIAL
Our statistical framework explains AF1’s potential as the update of a prior over dihedral angles with information on pairwise distances. For practical reasons, the application of PK by AF1 is heuristic in a few respects, which we clarify here.
First, the reference distribution over the coarse grained variable, $\pi(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi});\boldsymbol{\Gamma})$, is not directly obtained from the prior over the fine-grained variable, $\pi(\boldsymbol{\phi},\boldsymbol{\psi};\boldsymbol{\Theta})$. Rather, it is approximated by training a model that uses as input only the sequence length, $L$, and the positions of the Glycine residues, $\mathbb{I}_{\textrm{Gly}}(\boldsymbol{s})$.
Second, AF1 does not use proper PDFs over the distance matrix, $\boldsymbol{D}$. For $n$ points in $\mathbb{R}^{d}$, the n(n1)/2n(n-1)/2 pairwise distances cannot vary independently: they must satisfy geometric constraints to correspond to a valid 3D configuration. As discussed by Clifford and Green (1985), valid distances lie on a smaller manifold of dimension ndd(d+1)/2nd-d(d+1)/2. By modeling probability distribution over $\boldsymbol{D}$ as a product of independent marginals, AF1 defines an approximate distribution that is singular, i.e., it may place nonzero density on geometrically invalid configurations.